OpenStudy (anonymous):

find the value of k so that 2k+2, 5k-11, and 7k-13 is a geometric sequence.

mathslover (mathslover):

Can you find common ratio here?

OpenStudy (anonymous):

i think k is the common ratio

OpenStudy (cwrw238):

common ratio = 5k-11 / 2k+ 2 and this equals 7k-13 / 5k-11 solve this for k

OpenStudy (lgbasallote):

it didn't say those were consecutive terms though

mathslover (mathslover):

......common ratio in = (b)/ (a) where b and a are in geometric sequence

OpenStudy (anonymous):

@lgbasallote ..no

OpenStudy (lgbasallote):

@mathslover that only applies when b and a are consecutive

OpenStudy (lgbasallote):

@babybaby what do you mean no?

OpenStudy (anonymous):

it was not mentioned in my book that it is a geometric sequence. sorry

OpenStudy (cwrw238):

yes thats true of course - i assumed they were how else could we solve this

OpenStudy (lgbasallote):

...so it's not geometric sequence?

OpenStudy (cwrw238):

lol

OpenStudy (anonymous):

im sorry... what i mean is that it is not mentioned as consecutive

OpenStudy (lgbasallote):

this just got interesting

mathslover (mathslover):

Can we apply the sum of geometric sequence formula and then equating?

OpenStudy (cwrw238):

if they are consecutive equation reduces to 11k^2 - 98k + 147 = 0

OpenStudy (cwrw238):

i can't think of a way to do this unless they are consecutive terms

mathslover (mathslover):

........

OpenStudy (anonymous):

@ mathslover . yes

OpenStudy (anonymous):

lets try if it is a consecutive terms

OpenStudy (anonymous):

k=7

OpenStudy (anonymous):

@Algebraic! how did you answer it?

OpenStudy (anonymous):

same way as @cwrw238 evidently :)

OpenStudy (cwrw238):

solving the equation in last post gives k = 7

OpenStudy (anonymous):

set up 3 eqn.: a(r)^(n-1) = 2k+2 a(r)^n =5k -11 a*r^(n+1) = 7k-13

OpenStudy (anonymous):

how it was reduced?

OpenStudy (cwrw238):

also another value is 21/11

OpenStudy (anonymous):

start solving stuff in terms of k... eg. r= (5k-11)/(2x+2)

OpenStudy (anonymous):

@cwrw238 I saw that but I didn't check to make sure it worked... it does I guess..

OpenStudy (anonymous):

i knew it.

OpenStudy (anonymous):

thanks you guys. . .

OpenStudy (cwrw238):

7 gives a gp 16 , 24 , 36

OpenStudy (anonymous):

@cwrw238 aye

OpenStudy (cwrw238):

yw